Etchevers:Notebook/Genomics of hNCC/2008/08/09

From OpenWetWare
Jump to: navigation, search
C14.jpg Genomics of human neural crest cells Report.pngMain project page
Resultset previous.pngPrevious entry      Next entryResultset next.png

Discussions last week

Candice is undertaking a certain number of HIS.

  • The earliest one was on the C20 embryo she had sectioned, but did not work for LHX3/4 and not even very well for ISL1. Then again the + control did not work either.
  • Second HIS was with Emmanuelle for the candidate gene she discussed the other week) and that didn’t work either. Lots of blue background and not much signal. I hypothesized that under the current lab conditions of extreme heat, the pH of the Tris 9.5 when diluted to working 100mM (I think) might be too low for the NBT-BCIP step. Candice assured me that she is using fresh levamisole so that’s not it. It was 32ºC in the lab when we spoke, sometimes it’s 35º in the summer, as opposed to the optimal 20-22º.
  • We chose new slides for both the experiments; she will resynthesize some probes including Emmanuelle's gene (Tania) and LHX3 AS, but have new lot of ISL1 and will use VIM as + control also.
  • However, I took a look and the control (sense) slides are perfectly white. So it’s not a pH thing, and it’s not so uniformly blue as that. Bubble perhaps on the positive control slide (which definitely did not work – ISL1 near sections that had previously worked well)? Need to redo. Can then verify certain interesting structures eg. the developing cerebellum?

Finished up figure for Sarah/Laurence submission. Working on Puting's still (large figure, complex).

  • Alethea 11:07, 9 August 2008 (UTC):